# Ilias Bilionis

### v1

@@ Abstract
Accurate prediction of engine-out NOx is essential for meeting stringent emissions regulations and optimizing engine performance. Traditional approaches rely on models trained on data from a small number of engines, which can be insufficient in generalizing across an entire population of engines due to sensor biases and variations in input conditions. In real world applications, these models require tuning or calibration to maintain acceptable error tolerance when applied to other engines. This highlights the need for models that can adapt with minimal adjustments to accommodate engine-to-engine variability and sensor discrepancies. While previous studies have explored machine learning methods for predicting engine-out NOx, these approaches often fail to generalize reliably across different engines and operating environments. To address these issues, we propose a Bayesian calibration framework that combines Gaussian processes with approximate Bayesian computation to infer and correct sensor biases. Starting with a pre-trained model developed using nominal engine data, our method identifies engine specific sensor biases and recalibrates predictions accordingly. By incorporating these inferred biases, our approach generates posterior predictive distributions for engine-out NOx on unseen test data, achieving high accuracy without retraining the model. Our results demonstrate that this transferable modeling approach significantly improves the accuracy of predictions compared to conventional non-adaptive GP models, effectively addressing engine-to-engine variability and improving model generalizability.
@@ 1 Introduction
However, achieving high fidelity NOx prediction is technically challenging. NOx formation in engines is sensitive to many factors, such as combustion temperature, oxygen availability, fuel injection timing, and exhaust gas recirculation rates, which interact in a highly nonlinear fashion. Traditional NOx sensors are costly and operate only when exhaust temperatures are high enough, meaning there is no knowledge of engine-out NOx during cold exhaust conditions, such as engine startup. This has spurred interest in predictive models that estimate NOx emissions from readily available engine signals. Instead of substituting NOx sensors with a purely model based virtual sensor, model predictions can be used to improve the functionality of the physical sensors by providing a reference signal for consistency checks. For example, incorporating NOx prediction from the engine can improve the sensor signal integrity of physical sensors and help identify malfunctions or aging within the aftertreatment system [barbier2023predicting, ]. Furthermore, the estimation of real time NOx serves as a critical parameter for the vehicle’s on-board diagnostics system, allowing effective monitoring of engine states [walker2016future, ].
@@
In recent years, machine learning (ML) approaches have been extensively explored [aliramezani2022modeling, ] for predicting NOx emissions due to their ability to model complex, nonlinear relationships without requiring detailed physical insight into combustion phenomena. Artificial neural networks (ANNs) [fang2022artificial, ], support vector machines (SVMs) [aliramezani2020support, ], and deep neural networks (DNNs) [shin2020deep, ] have demonstrated high accuracy under both steady state and transient conditions. For instance,  [shin2020deep, ] designed a deep learning model to predict transient NOx emissions using steady state datasets, achieving notable accuracy by capturing input output relationships under dynamic conditions.  [fang2022artificial, ] used ANNs to predict transient NOx emissions from high speed direct injection diesel engines. Similarly,  [pillai2022modeling, ] developed DNN models for engine-out and tailpipe NOx prediction in heavy duty vehicles using onboard diagnostic data.  [aliramezani2020support, ] used SVMs to develop control-oriented NOx models, outperforming conventional regression algorithms in both transient and steady state conditions. Despite their success, these data-driven models often suffer from poor generalization when applied to unseen engine datasets and cannot respond to changes in the underlying physics of the system.
@@
To overcome the limitations of purely data-driven approaches, physics-based and semi empirical models incorporate underlying combustion principles to predict NOx emissions. For example,  [asprion2013fast, ] proposed a fast and accurate physics-based model for NOx. Their other work  [asprion2013optimisation, ] used simplified physical formulations to model NOx formation under real driving conditions. Their approach leveraged thermodynamic parameters such as combustion temperature and oxygen concentration to achieve computational efficiency. Similarly,  [mohammad2023physical, ] combined physical knowledge with dimensionality reduction to develop hybrid emission models, demonstrating that physically informed inputs improve NOx prediction accuracy.  [aithal2010modeling, ] used finite rate chemical kinetics to model NOx formation in diesel engines, highlighting the role of combustion processes. Another prominent study by  [lee2021real, ] developed real time NOx estimation models by embedding combustion parameters into a simplified physical framework. Although physics-based models are robust to variations in operating conditions, their reliance on empirical calibrations and in-cylinder measurements often limits their transferability across engines.
@@
Hybrid approaches combine the strengths of physics-based and ML methods to improve prediction accuracy and reduce computational costs. In these models, the combustion process is initially simulated using a straightforward physical model. The output of this physical simulation is then integrated with experimental data to identify the critical inputs required for training a data-driven methodology. This approach ensures that these models remain efficient and precise due to their reliance on data-driven techniques, while also maintaining adaptability to variations in the underlying physical phenomena, as it incorporates a physical combustion model [rezaei2020hybrid, ]. In  [mohammad2021hybrid, ], a hybrid framework was proposed where GT-SUITE, a 1D simulation tool, was coupled with SVMs and feedforward neural networks to predict NOx emissions. Their approach was evaluated on 772 steady‑state operating points from a 13‑liter heavy‑duty diesel engine, achieving an R2 of 0.99. However, its performance under transient conditions was not assessed.
@@
Despite advances in data-driven, physics-based, and hybrid techniques, the challenge of developing engine-to-engine transferable models remains largely unaddressed. Most of these models are engine specific and fail to generalize across engines of the same type due to differences such as manufacturing imperfections and mechanical wear. Ensuring engine-to-engine transferability with minimal retuning remains an open problem in the field. Therefore, developing transferable models that can accommodate this variablity is essential for improving scalability, reducing calibration costs, and allowing robust NOx prediction across multiple engines. There have been some attempts to improve transferability. For instance,  [shin2022designing, ] highlighted the domain mismatch problem between steady state and transient conditions, but did not explore engine-to-engine transferability. In another study  [shin2023task, ] used transfer learning in which knowledge from one task or dataset is transferred to speed up learning on a new task or engine. Their results showed that reusing a pretrained model can indeed save time and still achieve accuracy comparable to training a dedicated model for the new task or engine. However, this method required retraining for each new engine, a process that is computationally expensive and impossible for large scale deployment.
@@
Bayesian methods have shown potential to address model uncertainties and improve generalization. For example,  [cho2018structured, ] proposed a structured approach to uncertainty analysis of predictive models of engine-out NOx. Similarly, our previous work  [zinage2025causal, ] proposed a causal graph-enhanced Gaussian process (GP) regression to model engine-out NOx that uses a hybrid architecture using deep kernel learning [wilson2016deep, , zinage2024dkl, ]. However, this work has yet to be extended to tackle sensor biases and engine-to-engine variability effectively. To bridge this gap, this paper introduces a Bayesian framework that accounts for sensor biases while leveraging a pre-trained GP model. By including engine specific parameters, our approach achieves high predictive accuracy across different engines without retraining the GP model. This innovative solution not only improves engine-to-engine transferability but also quantifies uncertainty due to variability across different engines.
@@ 2 Engine-out NOx formation
The formation of nitrogen oxides (NOx), primarily comprising nitric oxide (NO) and nitrogen dioxide (NO2), in diesel compression ignition engines is a complex phenomenon governed by multiple chemical pathways. The well recognized pathways include thermal NOx, prompt NOx, and fuel bound NOx. Among these, the thermal NOx pathway, described by the extended Zeldovich mechanism, is the most significant contributor to engine-out NOx emissions in diesel engines.
@@ 3 Problem Statement
The core problem addressed in this work is to develop a Bayesian calibration framework for engine-out NOx models that (i) estimates sensor biases intrinsic to individual engines, (ii) adapts a pre-trained model to new engines with minimal adjustments, and (iii) produces reliable posterior predictive distributions (PPD) that quantifies uncertainty due to engine to engine variation without retraining.
@@ 4 Methodology
Formally, let $d_{\text{nc}}$ denote the number of non-control (measured) input variables. For engine $i$, we define the sensor bias vector as $b_{i}\in\mathbb{R}^{d_{\text{nc}}}$, assumed time invariant. To map these biases onto the full input vector, we introduce a fixed selection matrix $S\in\{0,1\}^{d\times d_{\text{nc}}}$, structured such that it selects only non-control inputs, leaving control inputs unchanged. The bias corrected input vector used for prediction is thus defined as:
@@
where the median is taken with respect to the GP predictive distribution and is used instead of the mean to increase robustness to skewness and outliers in the predictive distribution. The measurement model for engine-out NOx is then expressed as:
@@
(a) FTP cycle
(b) SET cycle
Figure 3: Variation in NOx measurements across different engines
@@ 5.1 Data Generation
The experimental datasets used in this study were provided by Cummins Inc. and consist of measurements from four multi-pulse fueling diesel compression ignition engines. For one of the engines, referred to as the “nominal engine”, both training and validation datasets were available. The training data were carefully designed to span the full range of operating conditions, ensuring the developed model is robust across all feasible engine states. For the remaining three engines, referred to as “sample engines”, only validation datasets were provided. These validation datasets were collected under conditions that mirrored the NOx levels and duty cycles of the validation data from the nominal engine. Throughout this investigation, the experimental measurements serve as ground truth and provide the benchmark for model performance evaluation.
@@ 5.2 Data Normalization
Standard normalization techniques such as min-max or standard scaler can inadequately handle extreme values, as these linear transformations remain significantly influenced by outliers. Consequently, we used the quantile transform normalization method [peterson2020ordered, ], which maps data onto a uniform or Gaussian distribution, thus providing robustness against outliers. By converting data points based on their relative ranking rather than their raw magnitude, quantile normalization effectively mitigates the disproportionate influence of extreme values, making it particularly suitable for complex, non-Gaussian emission datasets. All datasets used in this study were normalized using the quantile transform prior to GP model training on nominal engine data.
@@ 5.3 Metrics
Model evaluation was performed using multiple statistical metrics carefully chosen to capture diverse aspects of prediction accuracy especially under extreme and regulatory critical conditions:
@@
    effectively penalizes large prediction errors, aligning with the necessity of accurately predicting critical peak emission events.
  • Percentiles of Absolute Errors (90th, 95th, 98th): To assess performance during challenging conditions, we computed these higher order percentiles of the absolute prediction errors:
@@
    providing critical insights into the model’s reliability during infrequent but consequential emission scenarios.
Additionally, to qualitatively assess model performance, we used cumulative NOx plotted over time. This visualization highlights the overall alignment between predicted and observed NOx emissions, allowing intuitive identification of periods where the model consistently under or over estimates emissions.
@@ 6 Results
The GP model was implemented using GPyTorch [gardner2018gpytorch, ] with PyTorch as the backend. The loss function was defined as the negative of the exact marginal log-likelihood and the optimizer used is Adam [kingma2014adam, ] with tuned hyperparameters. To evaluate the ability of the models to accurately predict NOx emissions, we used several quantitative metrics, including the RMSE and the 90th, 95th, and 98th percentiles of absolute errors. The parameters used in the proposed approach are detailed in Table 1.
@@
Note that since a larger error was observed in NOx concentrations during the initial 80 seconds of the FTP cycle and the first 400 seconds of the SET cycle, possibly caused due to one engine warming up faster than the other, this data was excluded from analysis, and the subsequent 200 seconds (FTP) and 450 seconds (SET) (as shown in Fig. 3) were used to infer/estimate sensor biases. It is important to note that the objective is not to infer/estimate the sensor biases correctly but to find sensor biases that improve the predictions of engine-out NOx.
Once the sensor biases were inferred/estimated using the proposed approach, these biases were used to perform posterior predictive checks on the remaining data. Figures 4 and 5 show the PPD for the data used to infer sensor biases. This dataset, comprising 200 seconds from the FTP cycle and 450 seconds from the SET cycle for a specific sample engine, was used to select acceptable sensor biases based on the distance metric (i.e., KS statistic). We can see that the mean of the PPD for sample engines (engine 1, engine 2, and engine 3) closely matches the experimental data for the FTP cycle. However, for the SET cycle, the posterior mean exhibited relatively lower accuracy compared to the FTP cycle. This discrepancy aligns with the pretrained GP model’s reduced accuracy in predicting NOx concentrations for the SET cycle from the nominal engine, as shown in Fig. 6(b). Thus, such a pattern was expected for SET cycles from sample engines. The $95\%$ credible intervals, presented in Figures 4 and 5, reflect the uncertainty arising from the stochastic nature of inferred biases. The SET cycle displayed greater uncertainty, indicating a higher sensitivity of the biases to steady-state conditions.
@@
Fig. 7 presents the kernel density estimations (KDEs) of the prior and posterior samples of the sensor biases for engine-out NOx and non-control input variables for engine 1. We can see that the posterior concentrates around the values, which minimizes the distance between the output of the simulated model and the data samples that were used to infer these biases. Fig. 8 compares the predictions of the proposed approach with the conventional pre-trained GP model (i.e, GP (RBF) $W_{s}=5s$]) for the remaining data across both cycles. The median predictions of the proposed approach more closely align with the experimental data than those from the pretrained GP model which was trained on nominal engine data. Where discrepancies existed, the measurements largely fell within the uncertainty bands. Similar observations can be made for engines 2 and 3, as illustrated in Figures 9, 10, 11, and 12.
@@
To further validate the better performance of the proposed approach, Tables 2 and 3 present the RMSE and NOx absolute error percentiles for both models. The results demonstrate significant improvements in these metrics for the proposed approach compared to the GP (RBF) $W_{s}=5s$].

### v2

@@ Abstract
Accurate prediction of engine-out NOx is essential for meeting stringent emissions regulations and optimizing engine performance. Traditional approaches rely on models trained on data from a small number of engines, which can be insufficient in generalizing across an entire population of engines due to sensor biases and variations in input conditions. In real world applications, these models require tuning or calibration to maintain acceptable error tolerance when applied to other engines. This highlights the need for models that can adapt with minimal adjustments to accommodate engine-to-engine variability and sensor discrepancies. While previous studies have explored machine learning methods for predicting engine-out NOx, these approaches often fail to generalize reliably across different engines and operating environments. To address these issues, we propose a Bayesian calibration framework that combines Gaussian processes (GP) with approximate Bayesian computation to infer and correct sensor biases. Starting with a pre-trained model developed using nominal engine data, our method identifies engine specific sensor biases and recalibrates predictions accordingly. By incorporating these inferred biases, our approach generates posterior predictive distributions for engine-out NOx on unseen test data, achieving high accuracy without retraining the model. Our results demonstrate that this transferable modeling approach significantly improves the accuracy of predictions compared to conventional non-adaptive GP models, effectively addressing engine-to-engine variability and improving model generalizability.
@@ 1 Introduction
However, achieving high fidelity NOx prediction is technically challenging. NOx formation in engines is sensitive to many factors, such as combustion temperature, oxygen availability, fuel injection timing, and exhaust gas recirculation rates, which interact in a highly nonlinear fashion. Traditional NOx sensors are costly and operate only when exhaust temperatures are high enough, meaning there is no knowledge of engine-out NOx during cold exhaust conditions, such as engine startup. This limitation is particularly critical because a substantial fraction of the total cumulative tailpipe NOx is emitted during this cold-start and warm-up phase, before the aftertreatment system reaches its optimal operating temperature. This has spurred interest in predictive models that estimate engine-out NOx from readily available engine signals. Instead of substituting NOx sensors with a purely model based virtual sensor, model predictions can be used to improve the functionality of the physical sensors by providing a reference signal for consistency checks. For example, incorporating NOx prediction from the engine can improve the sensor signal integrity of physical sensors and help identify malfunctions or aging within the aftertreatment system [barbier2023predicting, ]. Furthermore, the estimation of real time NOx serves as a critical parameter for the vehicle’s on-board diagnostics system, allowing effective monitoring of engine states [walker2016future, ].
@@
In recent years, machine learning (ML) approaches have been extensively explored [aliramezani2022modeling, ] for predicting engine-out NOx due to their ability to model complex, nonlinear relationships without requiring detailed physical insight into combustion phenomena. Artificial neural networks (ANNs) [fang2022artificial, ], support vector machines (SVMs) [aliramezani2020support, ], and deep neural networks (DNNs) [shin2020deep, ] have demonstrated high accuracy under both steady state and transient conditions. For instance,  [shin2020deep, ] designed a deep learning model to predict transient engine-out NOx using steady state datasets, achieving notable accuracy by capturing input output relationships under dynamic conditions.  [fang2022artificial, ] used ANNs to predict transient engine-out NOx from high speed direct injection diesel engines. Similarly,  [pillai2022modeling, ] developed DNN models for engine-out and tailpipe NOx prediction in heavy duty vehicles using onboard diagnostic data.  [aliramezani2020support, ] used SVMs to develop control-oriented NOx models, outperforming conventional regression algorithms in both transient and steady state conditions. Despite their success, these data-driven models often suffer from poor generalization when applied to unseen engine datasets and cannot respond to changes in the underlying physics of the system.
@@
To overcome the limitations of purely data-driven approaches, physics-based and semi empirical models incorporate underlying combustion principles to predict engine-out NOx. For example,  [asprion2013fast, ] proposed a fast and accurate physics-based model for NOx. Their other work  [asprion2013optimisation, ] used simplified physical formulations to model NOx formation under real driving conditions. Their approach leveraged thermodynamic parameters such as combustion temperature and oxygen concentration to achieve computational efficiency. Similarly,  [mohammad2023physical, ] combined physical knowledge with dimensionality reduction to develop hybrid emission models, demonstrating that physically informed inputs improve NOx prediction accuracy.  [aithal2010modeling, ] used finite rate chemical kinetics to model NOx formation in diesel engines, highlighting the role of combustion processes. Another prominent study by  [lee2021real, ] developed real time NOx estimation models by embedding combustion parameters into a simplified physical framework. Although physics-based models are robust to variations in operating conditions, their reliance on empirical calibrations and in-cylinder measurements often limits their transferability across engines.
@@
Hybrid approaches combine the strengths of physics-based and ML methods to improve prediction accuracy and reduce computational costs. In these models, the combustion process is initially simulated using a straightforward physical model. The output of this physical simulation is then integrated with experimental data to identify the critical inputs required for training a data-driven methodology. This approach ensures that these models remain efficient and precise due to their reliance on data-driven techniques, while also maintaining adaptability to variations in the underlying physical phenomena, as it incorporates a physical combustion model [rezaei2020hybrid, ]. In  [mohammad2021hybrid, ], a hybrid framework was proposed where GT-SUITE, a 1D simulation tool, was coupled with SVMs and feedforward neural networks to predict engine-out NOx. Their approach was evaluated on 772 steady‑state operating points from a 13‑liter heavy‑duty diesel engine, achieving an $R^{2}$ of 0.99. However, its performance under transient conditions was not assessed.
Recent advances in physics informed neural networks [raissi2019physics, , nath2023physics, ] have further extended hybrid modeling paradigms by embedding governing physical laws directly into neural network training through physics based constraints in the loss function. Such approaches have demonstrated improved generalization and physical consistency in engine modeling tasks by incorporating thermodynamic relationships and combustion constraints into data driven architectures. While these methods improve model consistency with underlying physics, they are typically trained in a supervised manner on engine specific datasets, and their deployment across different engines may require additional adaptation or recalibration.
@@
Despite advances in data-driven, physics-based, and hybrid techniques, the challenge of developing engine-to-engine transferable models remains largely unaddressed. Most of these models are engine specific and fail to generalize across engines of the same type due to differences such as manufacturing imperfections and mechanical wear. Ensuring engine-to-engine transferability with minimal retuning remains an open problem in the field. Therefore, developing models that can accommodate this variability is essential for improving scalability, reducing calibration costs, and allowing robust NOx prediction across multiple engines. There have been some attempts to improve transferability. For instance,  [shin2022designing, ] highlighted the domain mismatch problem between steady state and transient conditions, but did not explore engine-to-engine transferability. In another study  [shin2023task, ] used transfer learning in which knowledge from one task or dataset is transferred to speed up learning on a new task or engine. Their results showed that reusing a pretrained model can indeed save time and still achieve accuracy comparable to training a dedicated model for the new task or engine. However, this method required retraining for each new engine, a process that is computationally expensive and impossible for large scale deployment.
@@
Bayesian methods have shown potential to address model uncertainties and improve generalization. For example,  [cho2018structured, ] proposed a structured approach to uncertainty analysis of predictive models of engine-out NOx. Similarly, our previous work  [zinage2025causal, ] proposed a causal graph-enhanced Gaussian process (GP) regression to model engine-out NOx. that uses a hybrid architecture using deep kernel learning [wilson2016deep, , zinage2024dkl, ]. However, this work has yet to be extended to tackle sensor biases and engine-to-engine variability effectively. To bridge this gap, this paper introduces a Bayesian framework that accounts for sensor biases while leveraging a pre-trained GP model. By including engine specific parameters, our approach achieves high predictive accuracy across different engines without retraining the GP model. This innovative solution not only improves engine-to-engine transferability but also quantifies uncertainty due to variability across different engines.
@@ 2 Engine-out NOx formation
The formation of nitrogen oxides (NOx) (Fig. 1), primarily comprising nitric oxide (NO) and nitrogen dioxide (NO2), in diesel compression ignition engines is a complex phenomenon governed by multiple chemical pathways. The well recognized pathways include thermal NOx, prompt NOx, and fuel bound NOx. Among these, the thermal NOx pathway, described by the extended Zeldovich mechanism, is the most significant contributor to engine-out NOx emissions in diesel engines.
@@
Figure 1: Formation of NOx in direct injected diesel compression ignition engines (Source: Figure adapted from khair2015nox )
@@ 3 Problem Statement
The core problem addressed in this work is to develop a Bayesian calibration framework for engine-out NOx models that simultaneously (i) estimates sensor biases intrinsic to individual engines, (ii) adapts a pre-trained model to new engine using only a small amount of engine-specific data and (iii) produces reliable posterior predictive distributions (PPDs) that quantify uncertainty due to engine-to-engine variability.
@@ 4 Methodology
Formally, let $d_{\text{nc}}$ denote the number of non-control (measured) input variables. For engine $i$, we define the sensor bias vector as $b_{i}\in\mathbb{R}^{d_{\text{nc}}}$, assumed time invariant. To map these biases onto the full input vector, we introduce a fixed selection matrix $S\in\{0,1\}^{d\times d_{\text{nc}}}$, structured such that it selects only non-control inputs, leaving control inputs unchanged. For instance, if the input vector $x$ is partitioned as $x=[u_{\text{control}},x_{\text{measured}}]^{T}$, then $S$ takes the block form $S=[0,I_{d_{\text{nc}}}]^{T}$. Here, $I_{d_{\text{nc}}}$ is the identity matrix corresponding to measured inputs (such as EGR system outlet temperature, intercooler outlet temperature, turbine inlet temperature) which are subject to sensor bias, and the zero matrix ensures that control inputs (such as VGT valve actuation, EGR valve actuation) remain uncorrected. The bias corrected input vector used for prediction is thus defined as:
@@
where the median is taken with respect to the GP predictive distribution and is used instead of the mean because the inverse transformation $Q^{-1}(\cdot)$ maps the GP outputs to a highly skewed, non-negative distribution in the physical domain.
(a) FTP cycle
(b) SET cycle
Figure 4: Variation in NOx measurements across different engines
In such heavy tailed distributions, the mean is often pulled toward extreme values, whereas the median remains a robust estimator of the central tendency and is invariant under the monotonic transformation $Q^{-1}$. The measurement model for engine-out NOx is then expressed as:
@@ 5.1 Data Generation
The experimental datasets used in this study were provided by Cummins Inc. and consist of measurements from four multi-pulse fueling diesel compression ignition engines. All engines belong to the Cummins B6.7L inline 6-cylinder diesel platform equipped with a high-pressure common rail fuel injection system, high-pressure EGR, and a VGT. The engines operate on ultra-low sulfur diesel fuel and are certified to EPA 2021 emissions standards. Key specifications of the engine are summarized in Table 1. The remaining sample engines share the same base architecture and emissions control configuration but exhibit engine-to-engine variability due to differences such as manufacturing imperfections and mechanical wear.
For one of the engines, referred to as the “nominal engine”, both training and validation datasets were available. The training data were carefully designed to span the full range of operating conditions, ensuring the developed model is robust across all feasible engine states. This nominal engine corresponds to the development engine in the test program with the most extensive instrumentation and complete coverage of the operating space. The remaining three engines, referred to as “sample engines”, are distinct production units of the same B6.7L platform, tested with the same fueling strategy, aftertreatment configuration, and fuel type. For the remaining three engines, referred to as “sample engines”, only validation datasets were provided. These validation datasets were collected under conditions that mirrored the NOx levels and duty cycles of the validation data from the nominal engine. Throughout this investigation, the experimental measurements serve as ground truth and provide the benchmark for model performance evaluation.
Table 1: Specifications of the engine
Specification Details
Engine type Cummins B6.7L CI engine
Horsepower 200-325 hp (149–242 kW)
Peak torque 520-750 lb-ft (705-1017 Nm)
Governed speed 2600 rpm
Clutch engagement torque 400 lb-ft (542 Nm)
Number of cylinders 6
Engine weight (dry) 1150 lb (522 kg)
Fuel system High pressure common rail
Turbocharger VGT
Emissions control High pressure EGR
Certification EPA 2021
@@ 5.2 Data Normalization
Furthermore, standard normalization techniques such as min-max or standard scaler can inadequately handle extreme values, as these linear transformations remain significantly influenced by outliers. Consequently, we used the quantile transform normalization method [peterson2020ordered, ], which maps data onto a uniform or Gaussian distribution, thus providing robustness against outliers. By converting data points based on their relative ranking rather than their raw magnitude, quantile normalization effectively mitigates the disproportionate influence of extreme values, making it particularly suitable for complex, non-Gaussian emission datasets. All datasets used in this study were normalized using the quantile transform prior to GP model training on nominal engine data.
@@ 5.3 Metrics
Model performance was evaluated using several statistical metrics carefully chosen to capture diverse aspects of prediction accuracy especially under extreme and regulatory critical conditions:
@@
    effectively penalizes large prediction errors, which is essential for accurately capturing peak emission events.
  • Percentiles of Absolute Errors (90th, 95th, 98th): To evaluate performance during challenging conditions, we computed these higher order percentiles of the absolute prediction errors:
@@
    providing essential information on model reliability in infrequent but impactful emission scenarios.
  • Coverage Probability of Credible Intervals: To assess uncertainty calibration, we computed the empirical coverage probability for the 95% posterior predictive credible intervals (CI):
    $\text{Coverage}_{95}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(y_{i}\in\left[\hat{y}_{i}^{2.5\%},\hat{y}_{i}^{97.5\%}\right]\right),$
    where $\hat{y}_{i}^{2.5\%}$ and $\hat{y}_{i}^{97.5\%}$ represents the lower and upper bounds of the interval, respectively. Well calibrated uncertainty estimates should yield coverage close to the nominal 95%.
Additionally, to qualitatively assess model performance, we used cumulative NOx plotted over time. This visualization highlights the overall alignment between predicted and observed engine-out NOx, allowing intuitive identification of periods where the model consistently under or over estimates emissions.
@@ 6 Results
The GP model was implemented using GPyTorch [gardner2018gpytorch, ] with PyTorch as the backend. The loss function was defined as the negative of the exact marginal log-likelihood and the optimizer used is Adam [kingma2014adam, ] with tuned hyperparameters. To evaluate the ability of the models to accurately predict engine-out NOx, we used several quantitative metrics, including the RMSE and the 90th, 95th, and 98th percentiles of absolute errors. The parameters used in the proposed approach are detailed in Table 2.
@@
All ABC experiments were executed on a desktop equipped with an Intel(R) Core(TM) i9-10900K CPU @ 3.70GHz and an NVIDIA RTX A4000 GPU with 16 GB of GDDR6 memory. Since the pretrained GP model remains fixed during calibration, the dominant computational cost arises from repeated GP median evaluations within the ABC simulation loop.
For each engine and cycle, the pilot phase used $N_{\text{pilot}}=1000$ samples and required approximately 20 seconds of wall-clock time. The main phase used $N_{\text{main}}=10000$ prior draws and required approximately 3–4 minutes per engine-cycle pair. The observed acceptance rate in the main phase was approximately $5\%$, consistent with the chosen quantile $\zeta=0.05$.
In total, the ABC procedure required approximately $N_{\text{pilot}}+N_{\text{main}}=11000$ parameter proposals per engine-cycle pair. Each proposal requires $T$ GP evaluations, resulting in approximately $2.2\times 10^{6}$ GP forward evaluations for the FTP cycle ($T=200$) and $4.95\times 10^{6}$ evaluations for the SET cycle ($T=450$). These evaluations are computationally efficient since they involve forward passes of a fixed GP model without gradient computation.
The observation noise variance $\sigma_{y}^{2}=0.01$ was selected based on the empirical variance of the normalized NOx residuals from the pretrained GP model on nominal validation data. Sensitivity analysis with $\sigma_{y}^{2}\in\{0.005,0.01,0.02\}$ resulted in negligible changes in posterior medians and predictive RMSE (less than 1%), confirming robustness to moderate misspecification. A sensitivity study was also conducted by varying $N_{\text{pilot}}\in\{500,1000,2000\}$ and $N_{\text{main}}\in\{5000,10000,20000\}$. We observed that reducing $N_{\text{pilot}}$ below 1000 resulted in slightly noisier tolerance estimates but negligible change in posterior medians. Increasing $N_{\text{main}}$ beyond 10000 produced minimal improvement in predictive metrics (less than 1% change in RMSE), indicating that the reported configuration in Table 2 achieves a stable tradeoff between computational cost and posterior accuracy.
@@
Note that since a larger error was observed in NOx concentrations during the initial 80 seconds of the FTP cycle and the first 400 seconds of the SET cycle, possibly caused due to one engine warming up faster than the other, this data was excluded from analysis, and the subsequent 200 seconds (FTP) and 450 seconds (SET) (as shown in Fig. 4) were used to infer/estimate sensor biases. To quantify the impact of excluding warm-up data, we performed an ablation study where biases were inferred using the dataset including the initial 80 seconds of FTP and 400 seconds of SET versus using only the stabilized portion. Table 3 presents the comparison. We can see that including warm-up data degraded prediction accuracy by 14–25% across all engines and cycles. This analysis revealed that during warm-up, several sensors exhibited non-stationary behavior inconsistent with the constant bias assumption, leading to biased parameter estimates that compromised predictions in the stabilized regime. These results justify our decision to exclude warm-up data for sensor bias calibration. It is important to note that the objective is not to infer/estimate the sensor biases correctly but to find sensor biases that improve the predictions of engine-out NOx.
Table 3: Impact of including warm-up data on prediction accuracy
Calibration Data RMSE
FTP cycle SET cycle
Engine 1
With warm-up (full) 94.3 128.4
Without warm-up (stabilized) 70.5 110.3
Improvement (%) 25.2% 14.1%
Engine 2
With warm-up (full) 108.7 135.2
Without warm-up (stabilized) 83.2 111.7
Improvement (%) 23.5% 17.4%
Engine 3
With warm-up (full) 90.4 120.1
Without warm-up (stabilized) 68.8 100.2
Improvement (%) 23.9% 16.6%
Once the sensor biases were inferred/estimated using the proposed approach, these biases were used to perform posterior predictive checks on the remaining data. Figures 5 and 6 show the PPD for the data used to infer sensor biases. This dataset, comprising 200 seconds from the FTP cycle and 450 seconds from the SET cycle for a specific sample engine, was used to select acceptable sensor biases based on the distance metric (i.e., KS statistic). We can see that the mean of the PPD for sample engines (engine 1, engine 2, and engine 3) closely matches the experimental data for the FTP cycle. However, for the SET cycle, the posterior mean exhibited relatively lower accuracy compared to the FTP cycle. This discrepancy aligns with the pretrained GP model’s reduced accuracy in predicting NOx concentrations for the SET cycle from the nominal engine, as shown in Fig. 7(b). Thus, such a pattern was expected for SET cycles from sample engines. The $95\%$ credible intervals, presented in Figures 5 and 6, reflect the uncertainty arising from the stochastic nature of inferred biases. The SET cycle displayed greater uncertainty, indicating a higher sensitivity of the biases to steady-state conditions. This behavior is consistent with the lower baseline accuracy of the pre-trained GP on steady-state data (Fig. 7(b)). Since the Bayesian calibration framework relies on the underlying model structure, the reduced fidelity in steady-state regimes naturally propagates into wider credible intervals in the PPD, ensuring that the model uncertainty accurately reflects the limitations of the base predictor.
@@
Fig. 8 presents the kernel density estimations (KDEs) of the prior and posterior samples of the sensor biases for engine-out NOx and non-control input variables for engine 1. We can see that the posterior concentrates around the values, which minimizes the distance between the output of the simulated model and the data samples that were used to infer these biases. Fig. 9 compares the predictions of the proposed approach with the conventional pre-trained GP model (i.e, GP (RBF) $W_{s}=5s$]) for the remaining data across both cycles. The median predictions of the proposed approach more closely align with the experimental data than those from the pretrained GP model which was trained on nominal engine data. Where discrepancies existed, the measurements largely fell within the uncertainty bands. Similar observations can be made for engines 2 and 3, as illustrated in Figures 10, 11, 12, and 13.
@@
To further validate the better performance of the proposed approach, Tables 4 and 5 present the RMSE and NOx absolute error percentiles for both models. The results demonstrate significant improvements in these metrics for the proposed approach compared to the GP (RBF) $W_{s}=5s$].
Table 6: Empirical coverage probability of 95% posterior predictive credible intervals
Engine FTP Cycle (%) SET Cycle (%)
Engine 1 96.2 87.1
Engine 2 96.1 86.3
Engine 3 87.5 88.6
Table 6 reports the empirical coverage probabilities of the 95% posterior predictive intervals. For the FTP cycle, coverage is close to the nominal 95% level across all engines, indicating well calibrated uncertainty. For the SET cycle, coverage remains above 85%, with slight undercoverage relative to nominal levels. This behavior is consistent with the reduced fidelity of the pretrained GP in steady state dominated regimes and confirms that the proposed framework provides reliable, though not overly conservative, uncertainty quantification.
Unlike the baseline pretrained GP model, which suffers from significant performance degradation due to engine-to-engine variability, the proposed framework offers a robust solution with minimal data requirements. Specifically, by using only a short segment of engine-specific data (approximately 200 to 450 seconds) for calibration, our method achieves a drastic reduction in predictive error, lowering the RMSE by up to 63% in transient cycles and 25% in steady state cycles compared to the non adaptive GP model. This capability not only ensures high fidelity predictions across different engines but also allow rapid, scalable deployment by avoiding the computational burden of retraining models for individual units in a fleet.